OpenStudy (anonymous):

Find the angle between v and w. Round your answer to one decimal place, if necessary. v = -5i + 7j, w = -6i - 4j

OpenStudy (anonymous):

that's what i have so far

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

4^2 is not 24

OpenStudy (anonymous):

wow. catch 1!

OpenStudy (amistre64):

the rest is good

OpenStudy (anonymous):

alright so 52 instead

OpenStudy (anonymous):

\[\sqrt{52}\]

OpenStudy (amistre64):

to save some wear and tear, (-5)^2 = (5)^2 so the signs arent that relevant to the length

OpenStudy (anonymous):

true, got it.

OpenStudy (amistre64):

and i like to stack the vectors that way the parts dont get 'confused' in the dot ai + bj xi + yj ------ ax + by

OpenStudy (amistre64):

you did fine other than 4^2 tho

OpenStudy (anonymous):

but to figure out the angle i would simplify the \[\theta = \cos^{-1} \frac{ 2 }{ \sqrt{74}\sqrt{52}}\]

OpenStudy (amistre64):

just stick it in the calculator ... no need to try to reduce and whatnot

OpenStudy (amistre64):

mkae sure your in degree mode if you want degrees

OpenStudy (amistre64):

sigh, its not 60

OpenStudy (amistre64):

6^2 + 4^2 is not going to be 60

OpenStudy (anonymous):

so 88.15 degrees

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

i had that in another tab before you corrected me the first time

OpenStudy (anonymous):

I swear i make more tiny mistakes than anything else.

OpenStudy (anonymous):

thank you again!

OpenStudy (amistre64):

the simple things really get you in the end :) i told my calculus teacher ... do you know how much algebra i had to forget for this test?

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

It even rings completely true!

OpenStudy (amistre64):

:) good luck